if f(x)=2x+4 and g(x)=x^2+1, find (f o g)(3)
Answer by Nate(3500)

Firstly, define [f o g].
f(x^2 + 1) = 2(x^2 + 1) + 4 = 2x^2 + 6
Now, plug in 3 for x.
2*9 + 6 = 24